A 54.6-kg skateboarder starts out with a speed of 1.65 m/s. He does 118 J of work on himself by pushing with his feet against the ground. In addition, friction does -269 J of work on him. In both cases, the forces doing the work are non-conservative. The final speed of the skateboarder is 7.00 m/s. (a) Calculate the change (PEf - PE0) in the gravitational potential energy. (b) How much has the vertical height of the skater changed? Give the absolute value.
Answer #1

work done by all forces on skateboarder = change in kinetic energy

work by himself +  work by friction + work by gravity = 1/2mVf^2 -1/2mVi^2 ( i,f are initial, final )

118 -269 + (PEf - PE0) = 1/2*54.6*(7^2-1.65^2) ( work by gravity = (PEf - PE0) )

=> (PEf - PE0) = 1414.37 J

=> mgh = (PEf - PE0)

=> h = 1414.37/(54.6*9.8) = 2.64 m